# On Secret-Message Transmission by Echoing Encrypted Probes

### v1

@@ Abstract
A scheme for secure communications, called “Secret-message Transmission by Echoing Encrypted Probes (STEEP)”, is revisited. STEEP is a round-trip scheme with a probing phase from one user to another and an echoing phase in the reverse direction. STEEP is broadly applicable to yield a positive secrecy rate in bits per channel use even if the receive channels at eavesdropper (Eve) are stronger than those between legitimate users in both directions. Differing from prior cases of STEEP, this paper presents STEEP in each of the following settings: using Gaussian channel probing (GCP) and Gaussian linear encryption (GLE) over MIMO channels; using phase-shift-keying (PSK) channel probing and PSK nonlinear encryption over SISO channels; and using GCP and GLE over multiple access channels. In each setting, Eve is assumed to have any number of antennas, and STEEP is shown to yield a positive secrecy rate subject to a sufficiently large power in the echoing phase relative to the power in the probing phase, as long as Eve’s receive channel in the probing phase is not noiseless. Also shown is that the GCP-GLE based STEEP for two users, subject to a form of large powers from the users, has its secrecy rate approaching the secret-key capacity based on GCP.
@@
Secure communications, information security, secret-key generation, secret-message transmission
@@ I Introduction
A notable exception is however the work in [22] where a two-way protocol for a binary symmetric channel system is proposed to achieve a positive secrecy rate even if Eve’s channel is stronger than users’. In fact, there is a general principle that predates and underpins this protocol, which consists of two integral steps:
@@
where $I(X;Y|Z)$ (for example) denotes the mutual information between $X$ and $Y$ conditional on $Z$. This is also known as Maurer’s bounds [7]. In some cases, the upper and lower bounds coincide, i.e., $C_{key}=I(X;Y)-I(Y;Z)=I(X;Y|Z)$ (e.g., see [11] for a Gaussian case), which is generally positive regardless of the WTC secrecy capacity $[I(X;Y)-I(X;Z)]^{+}$ from Alice to Bob. Here $x^{+}\doteq\max(x,0)$.
@@
The above two-step principle is also a foundation for a scheme called “Secret-message Transmission by Echoing Encrypted Probes (STEEP)” [10]. However, differing from [7], [2], [28] and [22], STEEP as shown in this paper allows the following extensions: $X$, $Y$ and $Z$ are allowed to be real, complex, vectors and/or matrices; the modulo sum $\oplus$ is allowed to be replaced by other suitable operations (examples will be shown); and the public channel from Bob to Alice and Eve may be replaced by any channels at the physical (or an upper) layer. Not necessarily all optimal in information theory, these extensions allow secure communications in a wider range of settings to be conducted rather simply with a guaranteed positive secrecy rate.
Figure 1: STEEP is a round-trip scheme with embedded secret message on returned (estimated) probes.
@@
The goal of this paper is to present STEEP in its latest forms. The primary contributions include novel insights into STEEP in three different settings. The first setting (or G-STEEP) uses Gaussian channel probing (GCP) and Gaussian linear encryption (GLE) over MIMO channels between two users, for which an achievable secrecy rate $C_{s,G}$ is derived and analyzed. In particular, $C_{s,G}$ is shown to converge to the secret-key capacity $C_{key}$ based on GCP over MIMO channels if the echoing power in G-STEEP dominates the probing power and both become large. The second setting (or P-STEEP) uses phase-shift-key (PSK) channel probing and PSK nonlinear encryption between two users, for which an achievable secrecy rate $C_{s,P}$ is also presented. The third setting (or M-STEEP) uses GCP and GLE over multiple access channels between an access point (AP) and multiple users all of whom apply the same probes from the AP. An achievable secrecy rate $C_{s,1}$ of M-STEEP from an arbitrary user to AP is shown to be a function that decreases gradually with some robustness (instead of abruptly) as the number $M$ of users increases. In each setting, the achievable secrecy rate of STEEP is shown to be positive subject to a sufficiently large power in the echoing phase. Note that “capacity” and “achievable rate” are treated interchangeably in this paper since each stated achievable rate is the maximum possible under stated conditions.
The paper is organized as follows. The physical-layer channel models of interest in this paper are described in section II. G-STEEP, P-STEEP and M-STEEP are presented and analyzed respectively in sections III, IV and V. Much of the technical proofs is relegated to the appendix. The paper is ended with additional comments and conclusion.
II Channel Models
Assume that the numbers of antennas on Alice, Bob and Eve are respectively $n_{A}$, $n_{B}$ and $n_{E}$. In the case of wireline communications, each antenna here corresponds to a transceiver.
When Alice transmits (within a coherence time $\mathcal{T}_{1}$) a sequence of random vectors $\sqrt{\frac{p_{A}}{n_{A}}}\mathbf{x}_{A}(k)\in\mathcal{C}^{n_{A}\times 1}$ of power $p_{A}$, we assume that Bob and Eve receive respectively
@@ II Channel Models
where all noises are normalized circular complex Gaussian noises, i.e., $\mathbf{w}_{B}(k)$ is $\mathcal{CN}(0,\mathbf{I}_{n_{B}})$ and $\mathbf{w}_{EA}(k)$ is $\mathcal{CN}(0,\mathbf{I}_{n_{E}})$. For notational simplicity, we will also use the scaled versions of $\mathbf{H}_{BA}$ and $\mathbf{H}_{EA}$, i.e., $\mathbf{H}_{BA}^{\prime}\doteq\sqrt{p_{A}/n_{A}}\mathbf{H}_{BA}\in\mathcal{C}^{n_{B}\times n_{A}}$ and $\mathbf{H}_{EA}^{\prime}\doteq\sqrt{p_{A}/n_{A}}\mathbf{H}_{EA}\in\mathcal{C}^{n_{E}\times n_{A}}$. Here $k$ is the sampling index.
Similarly, when Bob transmits (within a coherence time $\mathcal{T}_{2}$) a sequence of random vectors $\sqrt{\frac{p_{B}}{n_{B}}}\mathbf{x}_{B}(k)\in\mathcal{C}^{n_{B}\times 1}$ of power $p_{B}$, we assume that Alice and Eve receive respectively
@@
where the normalized noises $\mathbf{w}_{A}(k)$ and $\mathbf{w}_{EB}(k)$ are $\mathcal{CN}(\mathbf{0},\mathbf{I}_{n_{A}})$ and $\mathcal{CN}(\mathbf{0},\mathbf{I}_{n_{E}})$. We will also write $\mathbf{H}_{AB}^{\prime}=\sqrt{p_{B}/n_{B}}\mathbf{H}_{AB}\in\mathcal{C}^{n_{A}\times n_{B}}$ and $\mathbf{H}_{EB}^{\prime}=\sqrt{p_{B}/n_{B}}\mathbf{H}_{EB}\in\mathcal{C}^{n_{E}\times n_{B}}$.
Alice and Bob are half-duplex. Namely, $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ do not overlap. But $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ may or may not belong to a common coherence period.
Each receive channel state information (CSI) is assumed to be known to the corresponding receiver. If there is any CSI feedback between Alice and Bob, this CSI is also assumed to be known to Eve. In fact, all CSI in this paper is treated as known to Eve.
@@ III STEEP with Gaussian Channel Probing and Gaussian Linear Encryption (G-STEEP)
In this section, G-STEEP is presented, and an achievable secrecy rate $C_{s,G}$ of G-STEEP is also derived and discussed. Properties of $C_{s,G}$ subject to large powers are highlighted.
Given $n_{A}\geq n_{B}$, the probing phase should be from Alice to Bob in order to have the largest $C_{s,G}$. This is because the degree of freedom (DoF) of the secret-key capacity $C_{key}$ based on channel probing from a node with more antennas is larger than that from a node with less antennas. See [9] and [11]. Such a connection between $C_{s,G}$ and $C_{key}$ will also be shown.
III-A Phase 1 of G-STEEP
In phase 1, Alice applies Gaussian probing, i.e., she transmits a realization of the random probing vector $\sqrt{\frac{p_{A}}{n_{A}}}\mathbf{x}_{A}$ in each probing slot where $\mathbf{x}_{A}$ is assumed to be $\mathcal{CN}(0,\mathbf{I}_{n_{A}})$. The corresponding signal received by Bob is $\mathbf{y}_{B}$ in (4) (with “$(k)$” dropped). Let the (thin) SVD of $\mathbf{H}_{BA}$ be $\mathbf{U}_{BA}\boldsymbol{\Pi}_{BA}\mathbf{V}_{BA}^{H}=\sum_{i=1}^{n_{B}}\pi_{BA,i}\mathbf{u}_{BA,i}\mathbf{v}_{BA,i}^{H}$ where $\mathbf{U}_{BA}=[\mathbf{u}_{BA,1},\cdots,\mathbf{u}_{BA,n_{B}}]\in\mathcal{C}^{n_{B}\times n_{B}}$ is unitary and $\mathbf{V}_{BA}=[\mathbf{v}_{BA,1},\cdots,\mathbf{v}_{BA,n_{B}}]\in\mathcal{C}^{n_{A}\times n_{B}}$ is column-wise orthonormal. Then we can write
@@ III-A Phase 1 of G-STEEP
$\mathbf{R}_{\Delta\mathbf{p}}\doteq\mathbb{E}\{(\mathbf{\hat{p}}-\mathbf{p})(\mathbf{\hat{p}}-\mathbf{p}^{H})\}$
@@
III-B Phase 2 of G-STEEP
In phase 2, Bob applies Gaussian linear encryption, i.e., he transmits $\sqrt{\frac{p_{B}}{2n_{B}}}\mathbf{x}_{B}=\sqrt{\frac{p_{B}}{2n_{B}}}(\mathbf{\hat{p}}+\mathbf{s})$ where the secret message vector $\mathbf{s}$ is assumed to be $\mathcal{CN}(0,\mathbf{I}_{n_{B}})$. Here $p_{B}$ is the upper bound of the total transmit power from Bob. The corresponding signal received by Alice is
with $\mathbf{H}_{AB}^{\prime\prime}=\sqrt{1/2}\mathbf{H}_{AB}^{\prime}=\sqrt{p_{B}/(2n_{B})}\mathbf{H}_{AB}$. Assume that Alice and Eve both know the feedback of $\mathbf{V}_{BA}$ from Bob. Then, Alice also knows the effective probing vector $\mathbf{p}=\mathbf{V}_{BA}^{H}\mathbf{x}_{A}$.
Note that if $n_{A}=n_{B}=1$, then $\mathbf{V}_{BA}$ would reduce to one and hence the above mentioned feedback would not be needed.

### v2

@@ Abstract
A scheme for secure communications, called “Secret-message Transmission by Echoing Encrypted Probes (STEEP)”, is revisited. STEEP is a round-trip scheme with a probing phase from one user to another and an echoing phase in the reverse direction. STEEP is shown to be broadly applicable to yield a positive secrecy rate in bits per channel use even if the receive channels at eavesdropper (Eve) are stronger than those between legitimate users in both forward and reverse directions. This paper focuses on STEEP in the following settings: using Gaussian probing signal and Gaussian linear encryption over MIMO Gaussian channel (G-STEEP); using phase-shift-keying probing signal and a nonlinear encryption over SISO channel (P-STEEP); and a variation of G-STEEP for multiple access communication (M-STEEP). In each of the settings, Eve is assumed to have any given number of antennas, and STEEP is shown to yield a positive secrecy rate subject to a sufficiently large power in the echoing phase, as long as Eve’s receive channel in the probing phase is not noiseless. It is also shown that G-STEEP, subject to asymmetric large powers in forward and reverse directions, has its secrecy rate approaching the secret-key capacity based on Gaussian probing signal over MIMO Gaussian channel. STEEP does not require secure feedback channel, collaborative third party, in-band full-duplex or reciprocal channels between users, but only needs a design for echoing encrypted probes, asymmetric power allocation and/or collaborative round-trip coding.
@@
Secure communications, information security, secret-key generation, secret-message transmission.
@@ I Introduction
A notable exception is however the work in [22] where a two-way protocol using binary signalling over a Gaussian channel is proposed to achieve a positive secrecy rate even if Eve’s channel is stronger than users’. In fact, there is a general principle that predates and underpins this protocol, which consists of two integral steps:
@@
where $I(X;Y|Z)$ (for example) denotes the mutual information between $X$ and $Y$ conditional on $Z$. The left and right sides of (1) are known as Maurer’s lower and upper bounds [7]. In some cases (such as when $Y$ and $Z$ are independent of each other conditioned on $X$), the upper and lower bounds coincide, i.e., $C_{key}=I(X;Y)-I(Y;Z)=I(X;Y|Z)$, which is generally positive regardless of the WTC secrecy rate $[I(X;Y)-I(X;Z)]^{+}$ from Alice to Bob. Here $x^{+}\doteq\max(x,0)$. Note that $C_{key}$ is commonly referred to as a “capacity” despite the fact that it depends on the distributions of $X$, $Y$ and $Z$ that could be controllable in some applications.
@@
The above two-step principle is also a foundation for a scheme called “Secret-message Transmission by Echoing Encrypted Probes (STEEP)” [10]. However, differing from [7], [2], [28] and [22], STEEP as shown in this paper allows the following extensions: $X$, $Y$ and $Z$ are allowed to be any in the spaces of real and/or complex numbers and vectors and/or matrices; the modulo sum $\oplus$ is allowed to be replaced by other suitable operations (examples will be shown); and the public channel from Bob to Alice and Eve may be replaced by any channels at the physical (or an upper) layer. Not necessarily all optimal in information theory, these extensions allow secure communications in a wider range of settings to be conducted rather simply with a guaranteed positive secrecy rate.
Figure 1: STEEP is a round-trip scheme with embedded secret message on returned (estimated) probes. There is no requirement of secure feedback or in-band full duplex. The same probing signal transmitted/broadcasted by AP in phase 1 could be used by multiple users for orthogonal multiple access in phase 2.
@@
The goal of this paper is to present STEEP in its latest forms. The primary contributions include novel insights into STEEP in three different settings. The first setting (or G-STEEP) uses Gaussian probing signal (GPS) and Gaussian linear encryption (GLE) over MIMO channels between two users, for which an achievable secrecy rate $R_{s,G}$ is derived and analyzed. In particular, $R_{s,G}$ is shown to converge to the secret-key capacity $C_{key}$ based on GPS over MIMO channels if the echoing power in G-STEEP dominates the probing power and both become large. The second setting (or P-STEEP) uses phase-shift-key (PSK) probing signal and PSK nonlinear encryption between two users, for which an achievable secrecy rate $R_{s,P}$ is also presented. The third setting (or M-STEEP) uses GPS and GLE over multiple access channels between an access point (AP) and multiple users all of whom exploit the same probes from the AP. An achievable secrecy rate $\tilde{R}_{s,1}$ of M-STEEP from an arbitrary user to AP is shown to be a function that decreases gradually with some robustness (instead of abruptly) as the number $M$ of users increases. In each setting, the achievable secrecy rate of STEEP is shown to be positive subject to a sufficiently large power in the echoing phase, which includes the secrecy rate from a user to AP subject to exposure of messages from all other users.
The paper is organized as follows. The physical-layer channel models of interest in this paper are described in section II, which also highlights some prior results and the main problem of interest in this paper. G-STEEP, P-STEEP and M-STEEP are presented and analyzed respectively in sections III, IV and V. Much of the technical proofs is relegated to the appendix. The paper is ended with additional comments and conclusion.
II Channel Models, Prior Results, and The Problem
II-A Channel model
We will first consider a three-node network with two legitimate users (Alice and Bob) and an eavesdropper (Eve). The numbers of antennas on them are denoted respectively by $n_{A}$, $n_{B}$ and $n_{E}$. In the case of wireline communications, each antenna here corresponds to a transceiver.
When Alice transmits (within a coherence time $\mathcal{T}_{1}$) a sequence of random vectors $\sqrt{\frac{p_{A}}{n_{A}}}\mathbf{x}_{A}(k)\in\mathbb{C}^{n_{A}\times 1}$ of power $p_{A}$, we assume that Bob and Eve receive respectively
@@ II-A Channel model
where all noise entries are mutually independent with the normalized Gaussian distribution, i.e., $\mathbf{w}_{B}(k)$ is $\mathcal{CN}(0,\mathbf{I}_{n_{B}})$ and $\mathbf{w}_{EA}(k)$ is $\mathcal{CN}(0,\mathbf{I}_{n_{E}})$. For notational simplicity, we will also use the scaled versions of $\mathbf{H}_{BA}$ and $\mathbf{H}_{EA}$, i.e., $\mathbf{H}_{BA}^{\prime}\doteq\sqrt{p_{A}/n_{A}}\mathbf{H}_{BA}\in\mathbb{C}^{n_{B}\times n_{A}}$ and $\mathbf{H}_{EA}^{\prime}\doteq\sqrt{p_{A}/n_{A}}\mathbf{H}_{EA}\in\mathbb{C}^{n_{E}\times n_{A}}$. Here $k$ is the sampling index.
Similarly, when Bob transmits (within a coherence time $\mathcal{T}_{2}$) a sequence of random vectors $\sqrt{\frac{p_{B}}{n_{B}}}\mathbf{x}_{B}(k)\in\mathbb{C}^{n_{B}\times 1}$ of power $p_{B}$, we assume that Alice and Eve receive respectively
@@
where the normalized noises $\mathbf{w}_{A}(k)$ and $\mathbf{w}_{EB}(k)$ are $\mathcal{CN}(\mathbf{0},\mathbf{I}_{n_{A}})$ and $\mathcal{CN}(\mathbf{0},\mathbf{I}_{n_{E}})$. We will also write $\mathbf{H}_{AB}^{\prime}=\sqrt{p_{B}/n_{B}}\mathbf{H}_{AB}\in\mathbb{C}^{n_{A}\times n_{B}}$ and $\mathbf{H}_{EB}^{\prime}=\sqrt{p_{B}/n_{B}}\mathbf{H}_{EB}\in\mathbb{C}^{n_{E}\times n_{B}}$.
Alice and Bob are half-duplex (unless indicated otherwise). Namely, $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ do not overlap. But $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ may or may not belong to a common coherence period.
Every receive channel parameter is assumed to be known to the corresponding receiver. If there is any required feedback of channel parameters between users, these parameters are also assumed to be known to Eve. In fact, all channel parameters in this paper are treated as known to Eve.
@@
In section V, we will also consider an orthogonal multiple access problem where the access point (AP) has $n_{A}$ antennas and each of $M$ user equipment (UEs) has a single antenna. The channel parameters and noises are similarly defined.
II-B Some prior results and the problem
In the classic WTC scheme, the signals transmitted from Alice to Bob and Eve are not coordinated with any signals transmitted from Bob to Alice and Eve (even though both ends of a physical link are typically able to transmit). In this case, assuming all channel parameters are public, the secrecy capacity from Alice to Bob (in bits per complex channel use) is known [31, 32] to be
$C_{s,A\to B}=\max_{\mathbf{K}_{x},\texttt{Tr}\{\mathbf{K}_{x}\}\leq n_{A}}\log\frac{|\mathbf{I}_{n_{B}}+\frac{p_{A}}{n_{A}}\mathbf{H}_{B,A}\mathbf{K}_{x}\mathbf{H}_{B,A}^{H}|}{|\mathbf{I}_{n_{E}}+\frac{p_{A}}{n_{A}}\mathbf{H}_{E,A}\mathbf{K}_{x}\mathbf{H}_{E,A}^{H}|}$ (6)
where $\mathbf{K}_{x}=\mathbb{E}\{\mathbf{x}_{A}(k)\mathbf{x}_{A}^{H}(k)\}$. (The result for $C_{s,B\to A}$ would be obvious.) This secrecy capacity is achieved by a Gaussian codebook, i.e., $\mathbf{x}_{A}(k)$ follows the circular complex Gaussian distribution $\mathcal{CN}(\mathbf{0},\mathbf{K}_{x})$. Furthermore, $C_{s,A\to B}>0$ [31] if and only if (regardless of the positive power $p_{A}$ and, of course, $p_{B}$)
$\alpha\doteq\min_{\mathbf{v}\in\mathcal{C}^{n_{A}\times 1}}\frac{\|\mathbf{H}_{E,A}\mathbf{v}\|^{2}}{\|\mathbf{H}_{B,A}\mathbf{v}\|^{2}}<1.$ (7)
This means that Eve’s receive channel from Alice must be weaker than Bob’s receive channel from Alice in order for the classic WTC scheme to yield a positive secrecy rate from Alice to Bob. The above condition is however not always feasible. Specifically, when $n_{E}\geq n_{A}$, $\alpha$ is very likely larger than one in many practical situations especially where Eve is closer to Alice than Bob is.
Note that if $C_{s,A\to B}>0$, it is achieved by $\texttt{Tr}\{\mathbf{K}_{x}\}=n_{A}$ (using full power) [32]. But if $C_{s,A\to B}=0$, then it is obviously achieved by $\texttt{Tr}\{\mathbf{K}_{x}\}=0$ (using zero power), but not necessarily by $\texttt{Tr}\{\mathbf{K}_{x}\}=n_{A}$.
The main problem of interest in this paper is how to achieve a positive secrecy rate between two users (Alice and Bob), and between an access point and multiple user equipment, even if Eve’s receive channel is stronger than users’. In particular, we aim to present novel insights into STEEP and to reveal its ability to achieve a positive secrecy rate under virtually all channel conditions.
Subject to Gaussian distributed $\mathbf{x}_{A}(k)$ and any given $\mathbf{K}_{x}$, it is already established that there is a WTC coding scheme to yield a secrecy rate (see [2] among many sources):
$R_{s,A\to B}=\left(I(\mathbf{x}_{A}(k);\mathbf{y}_{B}(k))-I(\mathbf{x}_{A}(k);\mathbf{y}_{EA}(k))\right)^{+}$
$=\left(\log\frac{|\mathbf{I}_{n_{B}}+\frac{p_{A}}{n_{A}}\mathbf{H}_{B,A}\mathbf{K}_{x}\mathbf{H}_{B,A}^{H}|}{|\mathbf{I}_{n_{E}}+\frac{p_{A}}{n_{A}}\mathbf{H}_{E,A}\mathbf{K}_{x}\mathbf{H}_{E,A}^{H}|}\right)^{+}\leq C_{s,A\to B},$ (8)
where the equality holds when $n_{A}=1$. This result or its equivalent form will be used later for a number of channel conditions including virtual channel conditions. In particular, STEEP is a strategy that transforms a physical channel condition, even when (7) is not satisfied, into a virtual channel condition for which a positive secrecy rate can be ensured by a power control (i.e., collaboratively controlling $p_{A}$ and $p_{B}$).
@@ III STEEP with Gaussian Channel Probing and Gaussian Linear Encryption (G-STEEP)
In this section, G-STEEP is presented, and an achievable secrecy rate $R_{s,G}$ of G-STEEP is then derived and discussed. Properties of $R_{s,G}$ subject to large powers are highlighted.
III-A Description of G-STEEP
In phase 1 of G-STEEP, Alice applies Gaussian probing signal, i.e., she transmits a realization of the random probing vector $\sqrt{\frac{p_{A}}{n_{A}}}\mathbf{x}_{A}$ in each probing slot where $\mathbf{x}_{A}$ is assumed to be $\mathcal{CN}(0,\mathbf{I}_{n_{A}})$. The corresponding signal received by Bob is $\mathbf{y}_{B}$ in (4), i.e., after dropping “$k$”,
$\mathbf{y}_{B}=\mathbf{H}_{BA}^{\prime}\mathbf{x}_{A}+\mathbf{w}_{B}.$ (9)
Note that given $n_{A}\geq n_{B}$, the probing phase should be from Alice to Bob in order to have the largest $R_{s,G}$. This is because the degree of freedom (DoF) of the secret-key capacity $C_{key}$ based on channel probing from a node with more antennas is larger than that from a node with less antennas. See [9] and [11]. Such a connection between $R_{s,G}$ and $C_{key}$ will also be shown.
In phase 2 of G-STEEP, Bob transmits his estimated probing vector subject to a Gaussian linear encryption, i.e., he transmits $\sqrt{\frac{p_{B}}{2n_{B}}}\mathbf{x}_{B}=\sqrt{\frac{p_{B}}{2n_{B}}}(\mathbf{\hat{p}}+\mathbf{s})$ where $\mathbf{\hat{p}}$ is his estimate of the probing vector and $\mathbf{s}$ is a secret-message dependent vector and assumed to be $\mathcal{CN}(0,\mathbf{I}_{n_{B}})$. Here $p_{B}$ is the upper bound of the total transmit power from Bob. The corresponding signal received by Alice is
with $\mathbf{H}_{AB}^{\prime\prime}=\sqrt{1/2}\mathbf{H}_{AB}^{\prime}=\sqrt{p_{B}/(2n_{B})}\mathbf{H}_{AB}$.
The above protocol creates an effective wiretap channel (eWTC) system from $\mathbf{s}$ at Bob and the optimal estimate of $\mathbf{s}$ at Alice and the optimal estimate of $\mathbf{s}$ at Eve. We will show that this eWTC is in general in favor of the users.
III-B Analysis of the signal received by Bob in Phase 1
Let the (thin) SVD of $\mathbf{H}_{BA}$ be $\mathbf{U}_{BA}\boldsymbol{\Pi}_{BA}\mathbf{V}_{BA}^{H}=\sum_{i=1}^{n_{B}}\pi_{BA,i}\mathbf{u}_{BA,i}\mathbf{v}_{BA,i}^{H}$ where $\mathbf{U}_{BA}=[\mathbf{u}_{BA,1},\cdots,\mathbf{u}_{BA,n_{B}}]\in\mathbb{C}^{n_{B}\times n_{B}}$ is unitary and $\mathbf{V}_{BA}=[\mathbf{v}_{BA,1},\cdots,\mathbf{v}_{BA,n_{B}}]\in\mathbb{C}^{n_{A}\times n_{B}}$ is column-wise orthonormal. Then we can write
@@ III-B Analysis of the signal received by Bob in Phase 1
Assume that Alice and Eve both know the feedback of $\mathbf{V}_{BA}$ from Bob. Then, Alice also knows the effective probing vector $\mathbf{p}=\mathbf{V}_{BA}^{H}\mathbf{x}_{A}$. But if $n_{A}=n_{B}=1$, there is no need for feedback of $\mathbf{V}_{BA}$.
@@
$\mathbf{R}_{\Delta\mathbf{p}}\doteq\mathbb{E}\{(\mathbf{\hat{p}}-\mathbf{p})(\mathbf{\hat{p}}-\mathbf{p})^{H}\}=-\mathbb{E}\{(\mathbf{\hat{p}}-\mathbf{p})\mathbf{p}^{H}\}$
@@
$=\mathbf{I}_{n_{B}}-\boldsymbol{\Pi}_{BA}^{\prime 2}\left(\boldsymbol{\Pi}_{BA}^{\prime 2}+\mathbf{I}_{n_{B}}\right)^{-1}$
@@
III-C Analysis of the signal received by Alice in phase 2